# MCPZoo: A Large-Scale Dataset of Runnable Model Context Protocol Servers for AI Agent

### v2

@@ Abstract
Model Context Protocol (MCP) enables agents to interact with external tools, yet empirical research on MCP is hindered by the lack of large-scale, accessible datasets. We present MCPZoo, the largest and most comprehensive dataset of MCP servers collected from multiple public sources, comprising 95,142 servers. MCPZoo includes over ten thousand server instances that have been deployed and verified as runnable and interactable, supporting realistic experimentation beyond static analysis. The dataset provides unified metadata and access interfaces, enabling systematic exploration and interaction without manual deployment effort. MCPZoo is released as an open and accessible resource to support research on MCP-based security analysis.
@@ 1 Introduction
However, conducting large-scale empirical research on this growing ecosystem presents significant challenges. While recent studies have begun to explore the security risks and robustness of MCP integrations (e.g., prompt injection vulnerabilities in tool use), these efforts are often constrained by the difficulty of setting up testing environments. Unlike mobile apps which are packaged as standalone binaries, MCP servers are diverse software projects with complex, often conflicting dependencies. Consequently, existing research is frequently limited to small, manually curated datasets, leading to experiments that are difficult to reproduce or generalize.
@@
  • The largest and most comprehensive dataset of MCP servers currently available, with 95,142 MCP servers.
  • The first dataset to include more than ten thousand runnable and interactable server instances.
@@ 2.1 Construct MCP Zoo
Sources. We collected MCP server projects from seven public sources: MCP World [mcpworld], MCP Store [mcpstore], MCP Servers Repository [mcprepository], AIbase MCP [aibasemcp], Mcp.so [mcpso], MCP Market [mcpmarket], and Pulse MCP [pulsemcp]. Each source provides an independently curated directory of MCP servers, reflecting different community and platform perspectives.
@@ 3.1 Overview
MCPZoo is a large-scale, continuously growing dataset of Model Context Protocol (MCP) servers collected from multiple public sources. Up to December 17th, 2025, MCPZoo contains 95,142 distinct MCP servers, whose source code collectively amounts to 377 GB, among which 14,206 servers are verified to be runnable. Table I summarizes the composition of MCPZoo by source, reporting the number of collected servers from each of the platforms.
TABLE I: Statistics of MCP Servers from Seven Public Sources
@@
MCP World 31,048 32.63%
MCP Store 25,695 27.00%
MCP Repository 14,341 15.07%
AIbase MCP 11,120 11.69%
Mcp.so 6,772 7.12%
MCP Market 4,996 5.25%
Pulse MCP 1,170 1.23%
Total 95,142 100.0%
@@ 3.2 Visit MCPZoo
MCPZoo provides the capability to dynamically invoke MCP servers through remote connections. At present, 14,206 servers have been deployed, verified, and confirmed to be interactable in real execution environments. For each server, MCPZoo maintains a unified remote access configuration, which allows users to directly connect to and interact with MCP servers in a consistent manner. This design lowers the barrier to testing and experimentation. The MCPZoo website publicly exposes access interfaces for 15 runnable MCP servers for trial use. Researchers who require access to a larger set of interactable servers may contact us for extended access.
@@ 3.3 Access Conditions
4 Leverageing MCPZoo
MCPZoo can be used as a shared experimental foundation for research on real-world MCP deployments. For example, it enables agent benchmarking by providing diverse, realistic tool environments for evaluating agents’ planning and tool-use capabilities. It can also support security analysis at scale, such as systematically examining MCP servers for prompt injection or sandbox escape risks, extending prior studies on LLM and agent security to a protocol-level ecosystem. In addition, MCPZoo facilitates protocol evolution studies by allowing empirical analysis of MCP adoption and implementation patterns across a large set of real projects, in the spirit of prior large-scale software ecosystem measurements [androzoo].

### v3

@@ Abstract
Model Context Protocol (MCP) enables agents to interact with external tools, yet empirical research on MCP is hindered by the lack of large-scale, accessible datasets. We present MCPZoo, the largest and most comprehensive dataset of MCP servers collected from multiple public sources, comprising 129,059 servers (56,053 distinct). MCPZoo includes 16,356 server instances that have been deployed and verified as runnable and interactable, supporting realistic experimentation beyond static analysis. The dataset provides unified metadata and access interfaces, enabling systematic exploration and interaction without manual deployment effort. MCPZoo is released as an open and accessible resource to support research on MCP-based systems and security analysis.
@@ 1 Introduction
However, conducting large-scale empirical research on this growing ecosystem presents significant challenges. While recent studies have begun to explore the security risks and robustness of MCP integrations (e.g., prompt injection vulnerabilities in tool use), these efforts are often constrained by the difficulty of setting up testing environments. Unlike mobile apps, which are packaged as standalone binaries, MCP servers are diverse software projects with complex, often conflicting dependencies. Consequently, existing research is frequently limited to small, manually curated datasets, leading to experiments that are difficult to reproduce or generalize.
@@
  • The largest and most comprehensive dataset of MCP servers currently available, with 129,059 MCP servers.
  • The first dataset to include 16,356 runnable and interactable server instances.
@@ 2.1 Construct MCP Zoo
Sources. We collected MCP server projects from eight public sources: MCP Store [mcpstore], MCP World [mcpworld], MCP Market [mcpmarket], MCP Servers Repository [mcprepository], AIbase MCP [aibasemcp], Pulse MCP [pulsemcp], MCP.so [mcpso], and Smithery [smithery]. Each source provides an independently curated directory of MCP servers, reflecting different community and platform perspectives.
@@ 3.1 Overview
MCPZoo is a large-scale, continuously growing dataset of MCP servers collected from multiple public sources. Up to December 25th, 2025, MCPZoo contains 129,059 MCP servers, corresponding to 56,053 distinct servers after de-duplication. Among them, 16,356 servers are verified to be runnable. The collected source code amounts to 399 GB in total. Table I summarizes the composition of MCPZoo by source, reporting the number of collected servers from each of the platforms.
TABLE I: Statistics of MCP Servers from Multiple Public Sources (Dec. 2025)
@@
MCP Store 39,632 30.71%
MCP World 31,048 24.05%
MCP Market 16,105 12.48%
MCP Repository 14,341 11.11%
AIbase MCP 11,120 8.62%
Pulse MCP 6,884 5.33%
MCP.so 6,772 5.25%
Smithery 3,157 2.45%
Total 129,059 100.00%
Total (Distinct) 56,053
@@ 3.2 Visit MCPZoo
MCPZoo provides the capability to dynamically invoke MCP servers through remote connections. At present, 16,356 servers have been deployed, verified, and confirmed to be interactable in real execution environments. For each server, MCPZoo maintains a unified remote access configuration, which allows users to directly connect to and interact with MCP servers in a consistent manner. This design lowers the barrier to testing and experimentation. The MCPZoo website publicly exposes access interfaces for 15 runnable MCP servers for trial use. Researchers who require access to a larger set of interactable servers may contact us for extended access.
@@ 3.3 Access Conditions
4 Leveraging MCPZoo
MCPZoo can be used as a shared experimental foundation for research on real-world MCP deployments.
First, MCPZoo can enable principled agent benchmarking by providing diverse and realistic tool environments for evaluating agents’ capabilities. Benchmarking plays a critical role in emerging research areas [yang2025mcpsecbench, lian2025ase]. By supporting repeatable interaction with real MCP servers, MCPZoo makes it possible to construct authoritative, protocol-level benchmarks for MCP-based systems, complementing recent efforts that establish systematic evaluation frameworks for MCP security and AI-generated code at scale.
Second, MCPZoo can support large-scale security analysis from both attack and defense perspectives. Recent studies [guo2025systematicmcp, wang2025mcpguard] and industrial tools [mcp_scan2025, Tencent_AI-Infra-Guard_2025] have begun to identify various attack vectors against MCP systems, e.g., prompt injection, sandbox escape, and tool misuse. A large-scale, real-world dataset such as MCPZoo enables systematic examination of these vulnerabilities across diverse MCP implementations, moving beyond isolated case studies. At the same time, the availability of runnable servers provides a practical foundation for studying and validating defense mechanisms under realistic execution settings.
Finally, MCPZoo can facilitate ecosystem- and protocol-level evolution studies. For example, hou2025mcpreview systematically characterized security threats across the lifecycle of MCP. guo2025mcpecosystem conducted a large-scale empirical study of the MCP ecosystem from the perspectives of markets, servers, and clients, providing evidence-based insights into ecosystem-level security risks. By enabling empirical analysis of MCP adoption, implementation patterns, and cross-project diversity across a large collection of real-world MCP servers, MCPZoo offers a unique opportunity to study how the MCP ecosystem evolves in practice and to inform the future design of the protocol.